OpenStudy (anonymous):

What is the slope of the line that passes through the points (–5, 3) and (7, 5)? A. –1 B. 1/6 C. 1 D. 6

OpenStudy (breanna1999):

b

OpenStudy (anonymous):

why

OpenStudy (anonymous):

explain please

OpenStudy (breanna1999):

5-3/7-(-5)= 2/12= 1/6

OpenStudy (anonymous):

ok.thank you

OpenStudy (breanna1999):

welcome:)